OpenStudy (anonymous):

What is the value of n in the equation four left parenthesis three n plus two right parenthesis to the power of start fraction one over two end fraction end power minus five equals zero question mark

jimthompson5910 (jim_thompson5910):

can you draw out the problem? or post a screenshot?

OpenStudy (anonymous):

4(3n+2)1/2-5=0

jimthompson5910 (jim_thompson5910):

so this? \[\LARGE 4(3n+2)^{1/2} - 5 = 0\]

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

what's the first step? Any idea?

OpenStudy (anonymous):

no youre going to have to walk me through this one

jimthompson5910 (jim_thompson5910):

ok to make things simpler a bit, I'm going to let x = 3n+2 and replace the "3n+2" with x \[\LARGE 4(3n+2)^{1/2} - 5 = 0\] \[\LARGE 4x^{1/2} - 5 = 0\]

jimthompson5910 (jim_thompson5910):

what does that 1/2 power mean?

OpenStudy (anonymous):

not sure

jimthompson5910 (jim_thompson5910):

it means "square root", right?

jimthompson5910 (jim_thompson5910):

so \[\LARGE 4x^{1/2} - 5 = 0\] \[\LARGE 4\sqrt{x} - 5 = 0\]

jimthompson5910 (jim_thompson5910):

now try to isolate sqrt(x). Get sqrt(x) all by itself on one side

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you'll have to add 5 to one side

jimthompson5910 (jim_thompson5910):

yep, then what?

OpenStudy (anonymous):

not sure

jimthompson5910 (jim_thompson5910):

well you have this \[\LARGE 4\sqrt{x} = 5\]

jimthompson5910 (jim_thompson5910):

how would we move that 4 over

jimthompson5910 (jim_thompson5910):

hint: undo the "4 times"

OpenStudy (anonymous):

i reALLY DONT KNOW

jimthompson5910 (jim_thompson5910):

you divide both sides by 4 to undo the multiplication

jimthompson5910 (jim_thompson5910):

\[\LARGE 4\sqrt{x} = 5\] \[\LARGE \sqrt{x} = \frac{5}{4}\]

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

so whats the answer

jimthompson5910 (jim_thompson5910):

are you able to isolate x completely?

jimthompson5910 (jim_thompson5910):

square both sides

OpenStudy (anonymous):

i dont know

jimthompson5910 (jim_thompson5910):

give it a try

jimthompson5910 (jim_thompson5910):

don't be afraid to get it wrong

OpenStudy (anonymous):

i really dont

jimthompson5910 (jim_thompson5910):

5/4 times 5/4 = ???

OpenStudy (anonymous):

25/16

jimthompson5910 (jim_thompson5910):

remember we let x = 3n+2

jimthompson5910 (jim_thompson5910):

so if x = 25/16, then 3n+2 = 25/16

jimthompson5910 (jim_thompson5910):

isolate n in 3n+2 = 25/16

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

could you please hurry i have 3 more questions

jimthompson5910 (jim_thompson5910):

when you solve for n, what did you get?

OpenStudy (anonymous):

not sure

jimthompson5910 (jim_thompson5910):

what did you try so far

OpenStudy (anonymous):

i just want you to walk me through it bc i dont know how

jimthompson5910 (jim_thompson5910):

ok read this article and it might help out http://www.purplemath.com/modules/solvelin.htm let me know if it helps or not

OpenStudy (anonymous):

ok well could you please continue

jimthompson5910 (jim_thompson5910):

have a look at that article first

OpenStudy (anonymous):

i dont have time

jimthompson5910 (jim_thompson5910):

ok we can work on this when you get back

OpenStudy (anonymous):

no i need to know the answer